OpenStudy (anonymous):

@OrangeMaster The ratio of men to women working for a company 8 to 5 . If there are 70 women working for the company, what is the total number of employees? would i set up a proportion?

OpenStudy (anonymous):

M/W = 8/5 so 5 M = 8 W, M = (8/5) W also M+W = 70 (8/5) W + W = 70 solve for W and thus also get M

OpenStudy (anonymous):

i did it, and i got 112? did i do it wrong..

OpenStudy (anonymous):

but i did it a different way @douglaswinslowcooper

OpenStudy (anonymous):

No, I got it wrong. W = 70, not M+W= 70. Sorry. M = (8/5) 70 = 112. M+W = 70 + 112 = 182. Sorry to have led you astray.

OpenStudy (anonymous):

its okay!! thank you for the help i appreciate it

OpenStudy (anonymous):

You are quite welcome.